OpenStudy (richgall____):

Find the length of the unknown side. Round your answer to the nearest whole number. 7 inches 10 inches 25 inches 50 inches

OpenStudy (richgall____):

picture

OpenStudy (mrnood):

you can work this out 2 ways: since the 2 sides are equal you cqan see that the othe rside is the diagonal of a square:  The diagaonal of a square can be calculated by the multiplying the side by sqrt(2) The other way is a more general form of this and uses the pythage=oras equation a^2 + b^2 = c^2 in this case a & b = 5 so work out c from that equation